If the cube roots of unity are 1,w,w2 then the roots of the equation (x-13)+8=0 are..

Grade:12

1 Answers

Yash Baheti IIT Roorkee
9 years ago
Hi,

I think the question is (x-1)3 + 8 = 0
If that is the question, than you can write (x-1)3 = -8
or (x-1) = -2 and x = -1.

If its the one you have written than the answer is clrealy -7.